OpenStudy (henryarias14):

Determine whether the equation has a graph that is symmetric with respect to the y-axis, the x-axis, the origin, or none of these. 1) x2 - y2 = 4 2) y = -5x3 + 4x Instead of just the answer I'd like explanation too, please and thank you

OpenStudy (welshfella):

1) Note that when x = 2 y = 0 and when x = -2 y = 0

OpenStudy (welshfella):

what might this suggest?

OpenStudy (welshfella):

also there are 2 squared variables which suggests a type of conic section ( ellipse or hyperbole)

OpenStudy (henryarias14):

elipse? since they are even numbers?

OpenStudy (welshfella):

No - it has a negative sign so its a hyperbole and its symmetrical about which axis?

OpenStudy (henryarias14):

oh y-axis

OpenStudy (welshfella):

correct

OpenStudy (welshfella):

2) is a cubic polynomial nore that when x = 0 y = 0 so it passes through the origin

OpenStudy (henryarias14):

oh so the question is asking over what axis is the graph symetrical? what would be an example of a graph being symetrical on the origin? sorry i just have a hard time understanding what questions are asking

OpenStudy (henryarias14):

would an ellipse be symetrical on the origin?

OpenStudy (welshfella):

Sorry I lost the web x^2 + y^2 = 4 is a circle symmetrical about the origin Yes an ellipse could be symmetrical about the origin also

OpenStudy (welshfella):

No sorry - an ellipse might be symmetrical about the x or y axis,

OpenStudy (welshfella):

2) try solving -5x^2 + 4x = 0 to find where it cuts the x-axis

OpenStudy (welshfella):

* -5x^2 + 4x = 0

OpenStudy (welshfella):

gotta go - I forgot my soccer team are playing portugal in semifinal!

OpenStudy (henryarias14):

sorry i went to watch it too lol. good luck! Wales is playing resally good, its gonna be a nice game